A 1.55 gram sample pf pentanol, C5H11OH is burned and produced a temperature

increase of 55 0 C in 200 g of water. Calculate the heat of combustion of pentanol.


1
Expert's answer
2020-10-02T14:19:01-0400

At first, heat absorbed by water should be calculated.

Q=cm∆T, where c is a specific heat of water.

Q=4.184×200×55=46024(J).

It is equal (but with opposite sign) to heat released during combustion.

Q=-46024 J = -46.024 kJ.

To find molar heat of combustion this number should be divided into a mass of pentanol and multiplied by it's molar mass.

M=5×12+12×1+1×16=88(g/mol).

Q=-46.024×88/1.55=-2612.975(kJ).